OpenStudy (anonymous):

Write and simplify the equation you would use to solve the problem below. A farmer builds a rectangular pen with a perimeter of 116 feet. If the length is 5 feet more than the width, what is the width?

OpenStudy (e.cociuba):

Your equation will be 4w+10=116. Now solve it: 4w=116-10 w=106/4 w=26.5 ------------------- Now check you answer: 26.5*2 +(26.5+5)*2 =53+31.5*2 =53+63 =116